OpenStudy (anonymous):

What is the value of the expression d – 15 when d = –23? NEED HELP ASAP :(

OpenStudy (igreen):

d - 15 d = -23 -23 - 15 = -38 @peacepaw

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @peacepaw

OpenStudy (igreen):

If you need more help, just tag me by using the @ sign.